### v1

# Torus bifurcation of a dissipative time crystal

Abstract
Using a quantum gas setup consisting of a Bose-Einstein condensate strongly coupled to a high-finesse optical cavity by a transverse pump laser, we experimentally observe an instability of a dissipative continuous time crystal (CTC) towards a time crystalline state exhibiting two prominent oscillation frequencies. Applying a mean-field approximation model and a Floquet analysis, we theoretically confirm that this transition is a manifestation in a many-body system of a torus bifurcation between a limit cycle (LC) and a limit torus (LT). We theoretically illustrate the LC and LT attractors using the minimal model and experimentally reconstruct them using Takens’ embedding theorem applied to the non-destructively measured intracavity photon dynamics.
Introduction.— Critical transitions are ubiquitous in complex dynamical systems in nature. In condensed matter systems, a mean-field approximation of a many-body system can lead to a few-mode model, which maps certain aspects of a phase transition onto the mathematical framework of critical transitions and bifurcation theory typically applied to nonlinear systems [1]. For example, in dissipative systems, various phase transitions can be understood as critical transitions between two fixed points or a fixed point and a periodic solution [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. A physical system known to host a particularly rich variety of phase transitions is the quantum-gas cavity-QED platform [13, 14]. In the transversely pumped atom-cavity scenario, a pitchfork bifurcation from a trivial to a nontrivial fixed point is realised in the normal phase (NP) to superradiant phase (SP) transition as the strength of the light-matter interaction is increased [15, 16, 17, 18]. The emergence of a continuous time crystal (CTC), i.e., another many-body state, in that same scenario [19] can be understood as a Hopf bifurcation between the SP and a limit cycle (LC) as the light-matter interaction is further increased in a mean-field theory of the system [20, 21, 22, 23, 11]. Due to this correspondence between the many-body phase, the CTC, and its mean-field counterpart, the LC, we will refer to CTCs as LCs, hereafter. Within bifurcation theory, a LC can also become unstable as the system goes through a route to chaos [1]. An example is the period-doubling instability of a LC observed in a Rydberg atomic gas [24]. A less common type of instability of a LC, known to arise in classical systems [25, 26, 27, 28, 29], but yet to be experimentally demonstrated in quantum systems despite a theoretical proposal [30], is the Neimark-Sacker, secondary Hopf, or torus bifurcation. This bifurcation is characterised by a LC becoming unstable towards the formation of a limit torus (LT), which can be detected by observing the associated quasiperiodic dynamics.
In this work, we experimentally demonstrate that for strong light-matter interactions, the LCs previously observed in an atom-cavity system [19, 11], comprising a Bose-Einstein condensate (BEC) inside an recoil-resolved optical resonator [31, 32], becomes unstable, leading to a quasiperiodic dynamical state identified as a LT. The emergence of a LT corresponds to the appearance of a second spectral component in the Fourier spectrum of the intracavity photon dynamics, which is incommensurate to the response frequency of the LC. Supplementing our experimental results with numerical simulations, we infer from the behaviour of the Floquet multipliers of a minimal model of the system that the transition between a LC and LT can be understood as a Neimark-Sacker bifurcation, which constitutes its first observation in a quantum-coherent light-matter system.
Figure 1: (a) Schematic of the setup of a BEC transversally pumped with a standing wave potential. (b) Experimental phase diagram spanned by $\epsilon_{f}$ and $\delta_{\mathrm{eff}}$ showing distinct regimes: the (NP) normal phase, (SP) superradiant phase, (LC) limit cycle, (LT) limit torus, and (CH) chaos. (c), (d) Experimentally measured Fourier spectra of the intracavity photon number for (c) a LC and (d) a LT. (e) Fourier spectra for varying pump intensities at fixed detuning $\delta_{\mathrm{eff}}=-2\pi\times 2.6~{}\mathrm{kHz}$. The vertical lines denote the pump intensities corresponding to (c) and (d).
Experiment.— In the experiment, we prepare a BEC of 87Rb atoms with a particle number of around $N_{\mathrm{a}}\approx 4\times 10^{4}$. The BEC is strongly coupled to the fundamental mode of an optical high-finesse cavity by pumping with a retro-reflected light field along the direction perpendicular to the cavity axis as depicted in Fig. 1(a). We pump with a wavelength $\lambda_{\mathrm{p}}=791.59\,$nm, which is blue-detuned with respect to the $D_{1}$-line of 87Rb at 794.98794.98794.98\,794.98nm, and thus the pump field provides a repulsive standing wave potential for the atoms. The recoil frequency corresponding to this pump wavelength is $\omega_{\mathrm{rec}}=2\pi\times 3.7~{}\mathrm{kHz}$. The lattice depth produced by a single photon scattered into the cavity is $U_{\mathrm{0}}=2\pi\times 0.7\,$Hz. The bare detuning between the pump and cavity frequencies is $\delta_{c}$. When accounting for the the collective light shift, this gives rise to an effective detuning $\delta_{\mathrm{eff}}=\delta_{c}-N_{\mathrm{a}}U_{0}/2$. The cavity field decay rate is $\kappa=2\pi\times 3.2$ kHz. Thus, the system operates in the recoil-resolved regime $\kappa\sim\omega_{\mathrm{rec}}$, which means that the photon-mediated infinite-range interaction between the atoms is dynamical, and the cavity dynamics cannot be adiabatically eliminated, as for broad-bandwidth cavities with $\kappa\gg\omega_{\mathrm{rec}}$ [14].
In Fig. 1(b), the experimental phase diagram in the parameter space spanned by the final pump intensity $\epsilon_{\mathrm{f}}$ and the effective detuning $\delta_{\mathrm{eff}}$ is shown. The supplemental material contains details on the construction of the phase diagram and the criteria used to distinguish between the normal phase (NP), superradiant phase (SP), and chaotic response (CH) [33]. In the following, we will focus on the LC and LT. Later, we will demonstrate theoretically that the LT in this work appears as a torus in three-dimensional coordinate space. First, we will discuss the experimental signatures of the LC and LT based on the dynamics of the intracavity photon number $|a|^{2}$ used to construct the phase diagram in Fig. 1(b).
We analyse the normalised Fourier spectra of $|a|^{2}$, with some examples shown in Figs. 1(c) and 1(d) for a LC and LT, respectively. Both the LC and LT have well-defined oscillation frequencies appearing as peaks in the Fourier spectra of $|a|^{2}$. In a LC, there is one dominant response frequency, in addition to possible higher harmonics, as seen from the peak around $\sim$10 kHz in Fig. 1(c). In contrast, in a LT, a second prominent frequency peak is found, as shown in the subdominant peak around $\sim$7 kHz in Fig. 1(d). We emphasise that an incommensurate ratio between the primary and secondary peaks is indicative of a quasiperiodic dynamics, a defining feature of a LT. Note that a period-doubling bifurcation occurs when the ratio between the secondary and primary peaks is exactly $1/2$. In the Supplemental Material [33], we explain in detail how we experimentally classify LC and LT based on their unique signatures.
For fixed $\delta_{\mathrm{eff}}=-2\pi\times 2.6~{}\mathrm{kHz}$ and increasing pump intensity, we show in Fig. 1(e) that a LC becomes unstable and turns into a LT as a second peak in the spectra emerges. This is also supported by the island of LT found for higher values of $\epsilon_{\mathrm{f}}$ in the phase diagram Fig. 1(b). Therefore, we experimentally observe an instability of a LC towards a quasiperiodic dynamical state for increasingly strong light-matter interactions between the cavity photons and the atoms, the strength of which is effectively controlled by the pump intensity.
Multi-mode simulation.— To shed light on the nature of the transition between the LC and LT, we turn to theoretical descriptions of the transversely pumped atom-cavity system. We first employ a mean-field approach in a multimode expansion in momentum space. Later, we compare this approach to a few-mode mean-field approximation, which we discuss within a nonlinear dynamics methodology. For the multimode approach, we use a two-dimensional (2D) simplified model, considering only atomic momentum excitations along the pump and cavity directions together with the single cavity mode, as depicted in Fig. 2(a). We expand the atomic field operators using a plane-wave basis as done in Refs. [34, 35].
In our numerical simulations, we use the set of experimental parameters described earlier for an initial homogeneous BEC and a slightly occupied cavity mode $|a|^{2}(t=0)=10^{-3}$ to ensure that the system can be pushed out of the NP. We present in Fig. 2(b) the Fourier spectra for different pump intensities and fixed detuning $\delta_{\mathrm{eff}}=-2\pi\times 5~{}\mathrm{kHz}$ taken at $t\in[50,100]~{}\mathrm{ms}$; see Supplemental Material for the corresponding numerical phase diagram [33]. Consistent with the experimental results in Fig. 1(e), we also observe the appearance of side peaks that are incommensurate with the oscillation frequency of the previously stable LC before it turns into a LT for stronger $\epsilon$. Thus, our 2D model correctly predicts the instability observed in the experiment. We compare the predictions of the multimode model and the experimental observations to a few-mode model. In [11] we have demonstrated that it captures the Hopf bifurcation from a SP to a LC. As such, we will next focus on this minimal model to show that the LC-LT transition is a torus bifurcation and to show that a three-dimensional coordinate space is enough to reconstruct the LC and LT attractors.
Figure 2: Mean-field theory results. In (a), (b), a 2D model with 169 momentum modes and a single cavity mode is applied. The relevant modes associated with the momentum excitations in the two-mode approximation of the atomic sector for the three-mode model in (c) follow the same colour scheme. In (c)-(f) a minimal model with few coupled bosonic modes is used. (a), (c) Sketch of the theoretical models. (b), (d) Theoretical Fourier spectra for (b) $\delta_{\mathrm{eff}}=-2\pi\times 5~{}\mathrm{kHz}$ and (d) $-\delta_{\mathrm{eff}}=\omega_{p}=\kappa$. The maximum of the Fourier spectra is capped at $10^{-2}$ to emphasise the appearance of well-defined side peaks that are incommensurate to the dominant oscillation frequency of the LC. (e) Floquet multipliers $\mu$ for increasing $\lambda/\lambda_{\mathrm{SP}}\in[1.05,1.10]$ according to the minimal model. (f) Dynamics in a space spanned by $|a|$, $|b|$, and $|c|$ for $\lambda$ slightly smaller (dark) and larger (light) than the critical value for a LT to emerge.
Three-mode model.— The Hamiltonian for the minimal bosonic model reads [11]
$\hat{H}$ $=\omega_{p}\hat{a}^{\dagger}\hat{a}+\omega_{10}\hat{b}^{\dagger}\hat{b}+\omega_{20}\hat{c}^{\dagger}\hat{c}+\lambda\left(\hat{a}^{\dagger}+\hat{a}\right)\left(\hat{b}^{\dagger}+\hat{b}\right)$ (1)
$\qquad+\chi\hat{a}^{\dagger}\hat{a}\left(\hat{c}^{\dagger}+\hat{c}\right).$
In the context of the atom-cavity system, the operator $\hat{a}^{\dagger}$ creates a photon in the cavity mode. Excitations from the ground- state BEC mode $|0,0\rangle$ to the momentum modes |±k,±$|\pm\hbar k,\pm\hbar k\rangle$, crucial to the formation of a density wave pattern in the SP, are encapsulated in the bosonic creation operator $\hat{b}^{\dagger}$. The dynamical optical lattice formed along the cavity direction leads to $2\hbar k$-momentum excitations from $|0,0\rangle$ to $|0,\pm 2\hbar k\rangle$, which are represented by the creation operator $\hat{c}^{\dagger}$. This excitation channel couples to the intracavity intensity, and the coupling strength is controlled by the single photon lattice depth $U_{0}$ included in the prefactor $\chi$ in Eq. (1) [11]. We point out that this term is unique to our minimal model Eq. (1) and is absent in the usual treatment of a NP in the standard Dicke model in the thermodynamic limit [36]. The remaining parameters in Eq. (1) are the photon detuning $\omega_{p}=-\delta_{\mathrm{eff}}$, the standard light-matter coupling strength $\lambda$ proportional to the amplitude of the pump field, and the bare transition frequencies $\omega_{10}$ and $\omega_{20}$ corresponding to transitions from the ground state $|0,0\rangle$ to |±k,±$|\pm\hbar k,\pm\hbar k\rangle$ and $|0,\pm 2\hbar k\rangle$, respectively. A mean-field approximation of this model leads to a set of equations of motion (EOM) for $a=\langle\hat{a}\rangle$, $b=\langle\hat{b}\rangle$, and $c=\langle\hat{c}\rangle$ given by
$\frac{\mathrm{d}a}{\mathrm{d}t}$ $=-i\left(\omega_{p}-i\kappa+\chi\left(c+c^{*}\right)\right)a-i\lambda\left(b+b^{*}\right),$ (2)
$\frac{\mathrm{d}b}{\mathrm{d}t}$ $=-i\omega_{10}b-i\lambda\left(a+a^{*}\right),$
$\frac{\mathrm{d}c}{\mathrm{d}t}$ $=-i\omega_{20}c-i\chi a^{*}a~{}.$
We numerically propagate the above EOM for an initial configuration with slight occupation of the cavity mode and excited states, $a=b=c=10^{-3}$. The parameters are consistent with the experiment, i.e., $\omega_{10}=2\omega_{\mathrm{rec}}=\omega_{20}/2$, $\omega_{p}=-\delta_{\mathrm{eff}}$, and $\chi\approx 4\kappa$ [11]. We explore the dynamics for varying light-matter coupling strength in units of the critical value for the NP-SP transition, $\lambda_{\mathrm{cr}}=\sqrt{(\kappa^{2}+\omega_{p}^{2})\omega_{10}/4\omega_{p}}$. In Fig. 2(d), we present the Fourier spectra of $|a|^{2}$ for the minimal model taken at $\omega_{10}\,t\in[1,2]\times 10^{3}$. We point out the similarity in the general structure of the Fourier spectra of the 2D atom-cavity and minimal models, Figs. 1(b) and 1(d), which corroborates the applicability of our minimal model for gaining further insights about the LC-LT transition. We note that while the three-mode model in Eq. (2) can be further simplified to a model with two modes and a Kerr nonlinearity by an adiabatic elimination of the dynamics of the $c$-mode, such a model fails to capture the emergence of a Neimark-Sacker bifurcation and, instead, predicts a period-doubling bifurcation [33]. This highlights the crucial role played by the third atomic level for the existence of a Neimark-Sacker bifurcation in the system.
Floquet analysis.— Unlike in Ref. [11], which focused on the instability of the SP, a fixed point in the language of dynamical systems [1], we aim to identify in this work the type of bifurcation that leads to the instability of a LC, a periodic solution. As such, we employ a Floquet stability analysis [37, 38, 25, 39] of the LCs, which we briefly summarise as follows. First, we write Eq. (2) as $\partial_{t}\mathbf{X}=\mathbf{F}(\mathbf{X})$, where $\mathbf{X}=[a,a^{*},b,b^{*},c,c^{*}]^{\top}$, and numerically obtain the LC solution $\mathbf{X}_{\mathrm{LC}}(t)$ with period $T$. Next, we linearise Eq. (2) with respect to small perturbations around the stable LC solution, $\mathbf{X}_{\mathrm{LC}}(t)+\delta_{\mathbf{X}}$, which yields a linearised EOM $\partial_{t}\delta_{\mathbf{X}}=\mathbf{J}_{\mathrm{LC}}(t)\delta_{\mathbf{X}}$. Here, we have introduced the time-periodic Jacobian matrix $\mathbf{J}_{\mathrm{LC}}(t)$, which is the standard Jacobian matrix evaluated at the LC solution $\mathbf{J}_{\mathrm{LC}}(t)=\mathbf{J}_{0}\lvert_{\mathbf{X}_{\mathrm{LC}}}$, see Supplemental Material for an expression of the Jacobian matrix $\mathbf{J}_{0}$. We numerically obtain the solution of the linearised equation $\delta^{(1)}_{\mathbf{X}}(t)$ for an initial state that samples the vacuum fluctuations of the $a$-, $b$-, and $c$-modes, e.g., $a\in(\xi_{1}+i\xi_{2})/2\sqrt{N_{\mathrm{a}}}$, where $\xi_{i}$ is a random number drawn from a standard Gaussian distribution. We gather the solution as a column vector $[\delta^{(1)}_{\mathbf{X}}(t)]$ with $N=6$ rows corresponding to the dimension of the matrix $\mathbf{J}_{0}$ and the elements of which are the solutions $\delta^{(1)}$ evaluated at a specific time $t$. We obtain $N$ linearly independent solutions for $\delta_{\mathbf{X}}$ using a set of $N$ randomly picked initial states to construct the so-called fundamental matrix $\mathbf{\Delta}(t)=\biggl{[}\left[\delta^{(1)}_{\mathbf{X}}(t)\right],\dots,\left[\delta^{(N)}_{\mathbf{X}}(t)\right]\biggr{]}$. Finally, we construct the monodromy matrix $\mathbf{M}=\left(\mathbf{\Delta}(t)\right)^{-1}\mathbf{\Delta}(t+T)$, the eigenvalues of which are the Floquet multipliers $\mu_{i}$ with the property $\delta_{\mathbf{X}}(t+T)=\mu_{j}\delta_{\mathbf{X}}(t)$. Note that $\mu_{j}$ is, in general, a complex number as the monodromy matrix is typically non-Hermitian. For periodic solutions such as a LC, one Floquet multiplier $\mu_{0}$ is marginal or trivial as it is always equal to unity, $\mu_{0}=1$ [37, 38, 39]. This is because $\delta_{\mathbf{X}}=\dot{\mathbf{X}}_{\mathrm{LC}}$ is also a periodic solution of the linearised EOM and it has the same period $T$ as the LC. Excluding the marginal multiplier, a LC is stable if $|\mu_{j}|<1~{}\forall~{}j\neq 0$ [25]. On the other hand, if at least one of the Floquet multipliers has $|\mu_{j}|\geq 1$, then the LC is unstable.
Instability of a periodic solution is deduced if at least one nontrivial $\mu_{j}$ crosses the unit circle when the multipliers are plotted in a complex plane. For example, a period-doubling instability of a LC, as in Ref. [24], occurs when a purely real Floquet multiplier $\mu_{j}\to-1$ as the parameter for the transition is tuned [25, 29]. In fact, our simulations for a 1D atom-cavity system [33] predict a period-doubling instability of the LC, which suggests that an appropriate choice of transition frequencies in other setups [24, 40] could lead to realisation of other types of bifurcations. Here, for transition frequencies consistent with the appropriate excitation frequencies of the full atom-cavity setup, the behaviour of the Floquet multipliers for increasing light-matter coupling strength $\lambda$ is consistent with a Neimark-Sacker or torus bifurcation [25, 29]. This is evident from Fig. 2(e), in which we show the Floquet multipliers for different values of $\lambda$. Indeed, as $\lambda$ increases, a pair of $\mu_{j}$ eventually crosses the unit circle at an angle meaning $\mathrm{Im}(\mu_{j})>0$, consistent with a torus bifurcation [25, 29]. Hence, the LC-LT transition in the atom-cavity system in fact is a torus bifurcation.
Attractor reconstruction.— The attractor of a stable LC forms a closed loop in coordinate space while a LT forms a torus. According to our minimal model, a three-dimensional coordinate space spanned by $|a|$, $|b|$, and $|c|$ is sufficient to illustrate the LC and LT attractors in the atom-cavity system, as demonstrated in Fig. 2(f). This is especially insightful for the experimental reconstruction of the attractors in the atom-cavity setup, wherein the large number of atomic modes relevant in the dynamics makes it difficult to identify the minimum dimension of the coordinate space to show the attractor. Moreover, another limiting factor is that the destructive measurement of the dynamics of the atomic modes leads to a coarse scan of their time evolution. Nevertheless, attractors can be reconstructed using time delayed or lagged coordinates [41, 29, 40], which we can exploit together with the in situ monitoring of the cavity photon dynamics. Following Takens’ theorem, we delayed the single-shot intracavity photon occupation dynamics $|a|^{2}$ by $\tau=0.02$ ms and $2\tau$, as depicted in Figs. 3(a) and 3(b), and create a three-dimensional lagged coordinates for reconstructing the LC and LT attractors. The experimentally reconstructed attractors are shown in Fig. 3(c), wherein the trajectory for a LT is found to occupy a larger volume in the lagged-coordinate system than the LC. We attribute the apparent spiralling of the trajectory towards the centre of the LT attractor in Fig. 3(c) to atom losses, even though, in an idealized scenario the attractor should just remain a torus in the coordinate space, as highlighted by the theoretical results in Fig. 2(f).
Figure 3: Singlet trajectory measurement of the dynamics of the cavity photon occupation $|a|^{2}$ for (a) a LC and (b) a LT. The parameters for (a) and (b) are the same as those marked as ’X’ and ’O’ in Fig. 1(b). Each trajectory corresponds to the same data, albeit delayed by $\tau$ and $2\tau$, and shifted vertically by $40\times 10^{3}$ and $80\times 10^{3}$ from the original signal which is set to oscillate around zero, respectively. The time delay is $\tau=0.2$ ms. Vertical lines in (a) and (b) are guide to the eye to emphasize the time delay. (c) Experimental reconstruction of the LC and LT attractors using Takens’ embedding theorem using the time-delayed signals in (a) and (b).
Conclusion.— In conclusion, we have experimentally demonstrated a transition between two distinct time crystalline phases in a quantum gas setup. When approximated at the mean-field level, this transition exemplifies a Neimark-Sacker, or torus bifurcation, first predicted in nonlinear systems but now observed in a quantum many-body system. The torus bifurcation is observed when the CTC, initially defined by a single oscillation frequency, becomes unstable and begins to exhibit quasiperiodic oscillations characterised by two dominant frequencies. Theoretical analyses based on a minimal model and Floquet theory confirm that the quasiperiodic dynamical state is a LT and that the LC-LT transition can be understood as a torus bifurcation in the context of nonlinear dynamical systems. Our work opens the possibility to utilize the atom-cavity platform considered here or other similar driven-dissipative quantum systems operating in the few-body or mesoscopic regime as a natural testbed for exploring the quantum nature of bifurcations, limit cycles, and synchronization [42, 43, 44]. On the conceptual level, regarding phase transitions in and out of time crystalline states, we have provided support for the view that the criticality of time crystalline states is captured by bifurcation theory at the mean-field level.
Acknowledgements.
We thank J. Klinder and C. Georges for their support during the early stage of the project. This work was funded by the QuantERA II Programme that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No. 101017733, the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) “SFB-925” project 170620586, and the Cluster of Excellence “Advanced Imaging of Matter” (EXC 2056), Project No. 390715994. The project is co-financed by ERDF of the European Union and by ‘Fonds of the Hamburg Ministry of Science, Research, Equalities and Districts (BWFGB)’. J.G.C. and R.J.L.T. acknowledge support from the DOST-ASTI’s COARE high-performance computing facility. H.K. acknowledges funding by the state of North Rhine-Westphalia through the EIN Quantum NRW program. J.S. acknowledges support from the German Academic Scholarship Foundation.